OpenStudy (anonymous):

Type the ordered pair that is the solution to these equations. 2x - 3y = 7 -3x + y = 7

OpenStudy (anonymous):

multiply second eqn by3 and add to first eqn

OpenStudy (anonymous):

2x-3y=7 -9x+3y=21 so ans is -7x = 28 x = -4

OpenStudy (anonymous):

(-3 * -4) +y = 7 y = 7 - 12 y = -5

OpenStudy (anonymous):

So ordered pair is (-4,-5).